# Violations of the null convergence condition in kinematical transitions between singular and regular black holes, horizonless compact objects, and bounces

### v1

@@ Abstract:
How do regular black holes evade the Hawking–Penrose singularity theorems? Various models of stationary regular black holes, in particular those proposed by Bardeen, Dymnikova, Hayward, and close variants thereof, globally satisfy the null convergence condition (NCC). At first glance this might seem puzzling, as the NCC must generically be violated to avoid the focusing point implied by the singularity theorems. In fact, the extant singularity theorems depend on subtle global assumptions and do not provide information about where and when the singularity actually forms. In particular inner horizons are typically reached at finite affine parameter, before null geodesic focusing occurs, and the region inside the inner horizon is not itself a trapped region. Specifically, the Bardeen, Dymnikova, Hayward models of stationary regular black holes feature an inner Cauchy horizon which violates global hyperbolicity, hence violating one of the key assumptions of Penrose’s singularity theorem, and furthermore challenging their viability as long-living end-points of gravitational collapse. In contrast, during non-stationary processes describing kinematic transitions between standard singular black holes and regular black holes or horizonless compact objects, the inner horizon — when present — need not act as a Cauchy horizon. This raises the intriguing possibility that the NCC might instead be violated during intermediate stages of such transitions. Our detailed analysis confirms that NCC violations occur frequently during such kinematic transitions, even when the stationary end-point spacetimes respect the NCC. We also investigate analogous transitions toward black-bounce spacetimes and their horizonless compact counterparts, wormholes, where the NCC is always violated. These findings offer new insights into how regular black holes and related objects evade the constraints imposed by the standard singularity theorems.
Date: Sunday 2 February 2025 (NZT); -ed February 1, 2025.
@@ I Introduction
How do regular black holes (RBHs) evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems? This question is more subtle than one might suppose. Indeed, it is easy to check that the most common regular black holes (Bardeen, Dymnikova, Hayward, and close variants thereof Bardeen:1968bh ; Dymnikova:1992ux ; Hayward:2005gi ) globally satisfy the null convergence condition (NCC). [^1] More distantly related regular black holes, such as the Simpson–Visser black bounce Simpson:2018tsi (or the closely related Simpson–Visser naked traversable wormhole) do, however, globally violate the NCC. In contrast, the Minkowski-core regular black hole Simpson:2019mud ; Berry:2020ntz ; Simpson:2021dyo ; Simpson:2021zfl exhibits intermediate behaviour — the NCC is satisfied deep in the core and violated in the periphery. We focus on the geometric NCC, rather than the implicitly dynamical null energy condition (NEC) since the NEC makes dynamical assumptions regarding the applicability of the Einstein equations — we shall instead phrase the discussion in purely geometrical terms, eschewing the need for any particular choice of dynamics Carballo-Rubio:2019fnb . (For general discussions of energy conditions, in contrast to what can be done at a purely kinematic level, see, for instance, Curiel:2014zba ; Borde:1987qr ; Barcelo:2002bv ; Visser:1999de ; Martin-Moruno:2017exc ; Martin-Moruno:2013wfa ; Martin-Moruno:2013sfa ; Martin-Moruno:2015ena .) Since the most commonly occurring regular black holes satisfy the NCC, they must evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems in some other manner — the only remaining point of evasion being the hypothesized (global) causality condition.
Specifically, the presence of a static or stationary inner horizon, which is thereby a Cauchy horizon, violates global hyperbolicity. While improvements of the Penrose theorem, such as the Hawking–Penrose theorem, do not explicitly require the presence of a Cauchy hypersurface, still they implicitly assume some localized form of global hyperbolicity. (For instance, by postulating the absence of closed timelike curves; see e.g. Senovilla:2014gza ).
The above observations apply also to the much more familiar, and singular, Reissner–Nordström, Kerr, or Kerr–Newman spacetimes. It is perhaps a little sobering to realize that, due to the presence of Cauchy horizons, the standard Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems simply do not (directly) apply to any of the Reissner–Nordström, Kerr, or Kerr–Newman spacetimes. Certainly one can easily verify the presence of singularities in these spacetimes by direct computation, however, the presence of these singularities cannot be deduced from the standard Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems. (See in particular the closing comments in Senovilla:2014gza .)
@@
In non-stationary spacetimes the situation is much more complicated and interesting. If, for example, one constructs a kinematic model of a transition from/to a regular black hole to/from either a standard (singular) black hole, or to/from some horizonless compact object (HCO), then the evolving inner horizon (when present) need not be a Cauchy horizon. Similarly, in a dynamical gravitational collapse leading to the creation of a Kerr black hole only an inner horizon will need to be considered at finite times, at least before the formation of a time-like singularity, (which will then also imply the formation of a Cauchy horizon).
@@
If the inner horizon is not Cauchy, then there is no a priori reason to not apply the singularity theorems and hence the avoidance of a singularity implied by a late-time realization of one of the stationary regular black-hole geometries naturally leads to the conclusion that the NCC most probably has to be violated during the kinematic transition, though not at the stationary end-point. In this work, we shall verify that in several scenarios this is actually what happens: NCC violation can and does typically occur during the kinematic transition, even between end-point spacetimes that do themselves satisfy the NCC.
@@ IV Discussion
So how do regular black holes evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf ; Senovilla:2014gza ; Witten:2019qhl singularity theorems?
As pure mathematics, one can certainly write down many line elements that describe regular black holes all satisfying the NCC globally. This is certainly the case for the standard regular black hole spacetimes, [^6] If desired, these regular black hole “models” can sometimes be bootstrapped to “solutions” by reverse engineering a suitable matter Lagrangian — typically some variant on nonlinear electrodynamics. But this is not the main focus of the current article. such as Bardeen, Dymnikova, Hayward, and their close variants Bardeen:1968bh ; Dymnikova:1992ux ; Hayward:2005gi . [^7] More distantly related regular black holes, such as the black-bounce solutions (or the closely related naked traversable wormholes) do however violate the NCC globally. The intermediate “Minkowski core” regular black holes satisfy the NCC in the deep core, but violate it in the asymptotic region.
So, how do regular black holes actually form? One possibility is simply that the Hawking–Penrose singularity theorems cannot be applied to static or stationary regular black-hole gemoetries, as they are endowed, (in the same way as the singular Reissner–Nordström, Kerr, and Kerr–Newman spacetimes), with inner horizons which are also Cauchy horizons, and as such violate global hyperbolicity. This would, however, be a quite debatable resolution of the conundrum, as such geometries have to be considered as more mathematical abstractions, asymptotic states, rather than realistic models of the end points of a gravitational collapse.
(To put this in a somewhat blunter manner: The standard Hawking–Penrose singularity theorems are of limited utility in the presence of static or stationary inner horizons, including the Reissner–Nordström, Kerr, and Kerr–Newman spacetimes Senovilla:2014gza . Yes, there are finite affine-length null geodesics, but they accumulate on the Cauchy horizon — not on the singularity.)
On the other hand, in an evolving time-dependent situation, an inner horizon does not necessarily need to be a Cauchy horizon, and so it could be compatible with the assumption of global hyperbolicity. In such a situation, it seems one would be forced to violate the NCC in order to reach a singularity-free end point.
@@
We feel that this investigation serves to highlight more than ever the pressing need for improved singularity theorems, better-adapted to dealing with evolving inner and outer trapping horizons Hawking:2014tga ; Visser:2014zqa ; Ashtekar:2004cn . In particular, our investigations emphasize the need for theorems which are able to provide general conditions under which the formation of timelike singularities — shielded by dynamical inner horizons — is unavoidable. In conclusion, one might say that also in this situation, as has recently happened for mass inflation Carballo-Rubio:2024dca , regular black holes are forcing us to carefully reconsider the dynamical aspects of trapped regions and their horizons. As such, independently of their phenomenological interest, they ought to be considered crucial objects for current and future research in black hole physics (see e.g. Carballo-Rubio:2023mvr ; DiFilippo:2024spj ; Buoninfante:2024oxl ; Visser:2024zkx ; Carballo-Rubio:2025fnc ).

### v2

@@ Abstract:
How do regular black holes evade the Penrose singularity theorem? Various models of stationary regular black holes globally satisfy the null convergence condition (NCC). At first glance this might seem puzzling, as the NCC must generically be violated to avoid the focusing point implied by the Penrose theorem. In fact, the Penrose singularity theorem depends on subtle global assumptions and does not provide information about where and when the singularity actually forms. In particular inner horizons are typically reached at finite affine parameter, before null geodesic focusing occurs, and the region inside the inner horizon is not itself a trapped region. Specifically, the Bardeen, Dymnikova, Hayward models of stationary regular black holes feature an inner Cauchy horizon which violates global hyperbolicity, hence violating one of the key assumptions of Penrose’s singularity theorem, and furthermore challenging their viability as long-living end-points of gravitational collapse. In contrast, during non-stationary processes describing kinematic transitions between standard singular black holes and regular black holes or horizonless compact objects, the inner horizon — when present — need not act as a Cauchy horizon. This raises the intriguing possibility that the NCC might instead be violated during intermediate stages of such transitions. Our detailed analysis confirms that NCC violations occur frequently during such kinematic transitions, even when the stationary end-point spacetimes respect the NCC. We also investigate analogous transitions toward black-bounce spacetimes and their horizonless compact counterparts, wormholes, where the NCC is always violated. These findings offer new insights into how regular black holes and related objects evade the constraints imposed by the Penrose singularity theorem.
-ed February 12, 2025.
@@ I Introduction
How do regular black holes (RBHs) evade the Hawking–Penrose singularity theorems Penrose:1964wq ; Hawking:1970zqf (see also Senovilla:2014gza ; Hawking:1973uf )? This question is more subtle than one might suppose. In this paper we will focus on the Penrose (1965) singularity theorem Penrose:1964wq and treat the Hawking–Penrose (1970) singularity theorem Hawking:1970zqf in a companion paper ToAppear . The Penrose theorem establishes the null geodesic incompleteness of a spacetime $(\mathcal{M},g)$ under the following conditions:
  • i)
    null convergence condition (NCC): $R_{\mu\nu}\,\ell^{\mu}\ell^{\nu}\geq 0$ for all null vectors $\ell^{\mu}$,
  • ii)
    there is a non-compact Cauchy hypersurface $\Sigma$ in $\mathcal{M}$,
  • iii)
    there is a closed trapped surface $\mathcal{T}$ in $\mathcal{M}$
@@
It is easy to check that the most common regular black holes (Bardeen, Dymnikova, Hayward, and close variants thereof Bardeen:1968bh ; Dymnikova:1992ux ; Hayward:2005gi ) globally satisfy the NCC. [^1] More distantly related regular black holes, such as the Simpson–Visser black bounce Simpson:2018tsi (or the closely related Simpson–Visser naked traversable wormhole) do, however, globally violate the NCC. In contrast, the Minkowski-core regular black hole Simpson:2019mud ; Berry:2020ntz ; Simpson:2021dyo ; Simpson:2021zfl exhibits intermediate behaviour — the NCC is satisfied deep in the core and violated in the periphery. We focus on the geometric NCC, rather than the implicitly dynamical null energy condition (NEC) since the NEC makes dynamical assumptions regarding the applicability of the Einstein equations — we shall instead phrase the discussion in purely geometrical terms, eschewing the need for any particular choice of dynamics Carballo-Rubio:2019fnb . (For general discussions of energy conditions, in contrast to what can be done at a purely kinematic level, see, for instance, Curiel:2014zba ; Borde:1987qr ; Barcelo:2002bv ; Visser:1999de ; Martin-Moruno:2017exc ; Martin-Moruno:2013wfa ; Martin-Moruno:2013sfa ; Martin-Moruno:2015ena .) Since the most commonly occurring regular black holes satisfy the NCC, they must evade the Penrose singularity theorem in some other manner — the only remaining point of evasion being the hypothesized (global) causality condition which postulates the existence of a non-compact Cauchy hypersurface, tantamount to demanding global hyperbolicity Specifically, the presence of a static or stationary inner horizon, which is thereby a Cauchy horizon, violates global hyperbolicity.
We emphasize that the Hawking–Penrose (1970) singularity theorem Hawking:1970zqf (see also Senovilla:2014gza ; Hawking:1973uf ) explicity removes the assumption of existence of a Cauchy hypersurface by instead demanding the generic condition for the curvature tensor and the absence of closed timelike curves, but most importantly assumes the timelike convergence condition (TCC) in addition to the NCC. Concerning the Hawking–Penrose theorem (1970), it is this last condition which is violated by the standard de Sitter core regular black holes.
The above observations apply also to the much more familiar and singular Reissner–Nordström spacetime. It is perhaps a little sobering to realize that, due to the presence of a Cauchy horizon, the standard Penrose singularity theorem does not apply. The Hawking–Penrose theorem, however, does establish the geodesic incompleteness of the Reissner–Nordström spacetime. Its regular counterpart, the Bardeen black hole Bardeen:1968bh , evades this theorem purely by the fact that it violates the TCC. We analyze such TCC violations in a companion paper ToAppear .
@@
Regarding the Penrose (1965) theorem, in non-stationary spacetimes the situation is much more complicated and interesting. If, for example, one constructs a kinematic model of a transition from/to a regular black hole to/from either a standard (singular) black hole, or to/from some horizonless compact object (HCO), then the evolving inner horizon (when present) need not be a Cauchy horizon. Similarly, in a dynamical gravitational collapse leading to the creation of a Kerr black hole only an inner horizon will need to be considered at finite times, at least before the formation of a time-like singularity, (which will then also imply the formation of a Cauchy horizon).
@@
If the inner horizon is not Cauchy, then there is no a priori reason to not apply the Penrose singularity theorem and hence the avoidance of a singularity implied by a late-time realization of one of the stationary regular black-hole geometries naturally leads to the conclusion that the NCC most probably has to be violated during the kinematic transition, though not at the stationary end-point. In this work, we shall verify that in several scenarios this is actually what happens: NCC violation can and does typically occur during the kinematic transition, even between end-point spacetimes that do themselves satisfy the NCC.
@@ IV Discussion
So how do regular black holes evade the Hawking–Penrose Penrose:1964wq ; Hawking:1970zqf singularity theorems (see also Senovilla:2014gza ; Hawking:1973uf )? In this work we have addressed different aspects of this question focussing on the Penrose (1965) theorem Penrose:1964wq , whereas the Hawking–Penrose (1970) theorem Hawking:1970zqf will be discussed elsewhere ToAppear .
As pointed out, for various stationary regular black hole models, it is the inner horizon, which is also a Cauchy horizon, that violates one of the assumptions of the Penrose singularity theorem, despite these models satisfying the NCC globally.
So, how do regular black holes actually form? One possibility is simply that the Penrose singularity theorem cannot be applied to static or stationary regular black-hole gemoetries, as they are endowed, (in the same way as the singular Reissner–Nordström spacetime), with inner horizons which are also Cauchy horizons, and as such violate global hyperbolicity. This would, however, be a quite debatable resolution of the conundrum, as such geometries have to be considered as more mathematical abstractions, asymptotic states, rather than realistic models of the end points of a gravitational collapse. Instead, it is reasonable to expect that the Hawking–Penrose (1970) singularity theorem provides insights into this question. The latter assumes the TCC, in addition to the NCC, but relaxes the assumption of existence of a non-compact Cauchy hypersurface. In particular, this theorem establishes the geodesic incompleteness of the Reissner–Nordström spacetime and exemplifies how de Sitter core regular black holes avoid a singularity by violating the TCC.
Regarding the Penrose singularity theorem, on the other hand, in an evolving time-dependent situation, an inner horizon does not necessarily need to be a Cauchy horizon, and so it could be compatible with the assumption of global hyperbolicity. In such a situation, it seems one would be forced to violate the NCC in order to reach a singularity-free end point.
@@
We feel that this investigation serves to highlight more than ever the pressing need for improved singularity theorems better-adapted to dealing with evolving inner and outer trapping horizons Hawking:2014tga ; Visser:2014zqa ; Ashtekar:2004cn , as well as an improved understanding of existent singularity theorems and their global and geometric convergence conditions. In particular, our investigations emphasize the need for an improved understanding of general conditions under which the formation of timelike singularities — shielded by dynamical inner horizons — is unavoidable. In conclusion, one might say that also in this situation, as has recently happened for mass inflation Carballo-Rubio:2024dca , regular black holes are forcing us to carefully reconsider the dynamical aspects of trapped regions and their horizons. As such, independently of their phenomenological interest, they ought to be considered crucial objects for current and future research in black hole physics (see e.g. Carballo-Rubio:2023mvr ; DiFilippo:2024spj ; Buoninfante:2024oxl ; Visser:2024zkx ; Carballo-Rubio:2025fnc ).